The area of a square is 5 more times than its perimeter. Find the length of a side.
Answer by nerdybill(7384)

The area of a square is 5 times than its perimeter. Find the length of a side.
.
Let x = length of one side of square
then
"area of square" = "5 more times than its perimeter"
x^2 = 5(4x)
x^2 = 20x
x^2 - 20x = 0
x(x - 20) = 0
x = {0, 20}
.
length of side must be 20